OpenStudy (anonymous):

Another plan to secure the roller coaster involves placing two concrete struts on either side of the center of the leg of the roller coaster to add reinforcement against southerly winds in the region. Again, using the center of the half-circle as the origin, the struts are modeled by the equations and . A vertical reinforcement beam will extend from one strut to the other when the two cables are 2 feet apart.

OpenStudy (anonymous):

Equations are \[y=\sqrt{x+8} and y=\sqrt{x-4}\]

OpenStudy (anonymous):

again, it can be found here

OpenStudy (radar):

Those equations are not consistent, or at least I can not find a solution.

OpenStudy (anonymous):

on the assignment itself, model 3 number 8.

OpenStudy (anonymous):

I'm not sure if i put the information out right.

OpenStudy (radar):

I am unable to display the diagram on my computer.

OpenStudy (radar):

For some reason my open office has crashed. I need to log off.

OpenStudy (anonymous):

Can i take a screenshot?

OpenStudy (radar):

I think that might work. I am using a Mac

OpenStudy (anonymous):

http://imgur.com/7dbjeoO

OpenStudy (radar):

O.K. I have it, and now I am trying to figure what the problem consist and form the needed equations

OpenStudy (anonymous):

ok, thank you!

OpenStudy (radar):

I just can't figure it out, I think they are wanting x when the vertical (y) difference is 2 ft. If I am correct (which I don't feel to positive about) then we would have this situation:\[\sqrt{x+8}-\sqrt{x-4}= 2\]becoming\[\sqrt{x+8}= 2 + \sqrt{x-4} \]Squaring both sides we then get:\[8=4\sqrt{x-4}\]dividing both sides by 4\[2=\sqrt{x-4}\]squaring both sides getting: 4 = x - 4 or x = 8. Here I am treading in unfamiliar territory but this is all I can think of at the moment. I am calling it quits for tonight. Good luck with it.

OpenStudy (anonymous):

Thank you very much, you were a great help!